A boxing ring is in the shape of a square, 20 ft on each. How far apart are the fighters when they are in opposite corners of the ring?
Answer by checkley77(12844)

You are looking for the hypotenuse of a right triangle with equal sides of 20 ft.
20^2+20^2=d^2
400+400=d^2
d^2=800
d=sqrt800
d=28.28 ft. apart when in the opposite corners.